# Synthetic Data Generation and Differential Privacy using Tensor Networks’ Matrix Product States (MPS)

### v1

@@ Abstract
Synthetic data generation is a key technique in modern artificial intelligence, addressing data scarcity, privacy constraints, and the need for diverse datasets in training robust models. In this work, we propose a method for generating privacy preserving high-quality synthetic tabular data using Tensor Networks, specifically Matrix Product States (MPS) [1].We benchmark the MPS-based generative model against state-of-the-art models such as CTGAN [2], VAE, and PrivBayes [3], focusing on both fidelity and privacy-preserving capabilities. To ensure differential privacy (DP), we integrate noise injection and gradient clipping during training, enabling privacy guarantees via Rényi Differential Privacy accounting [4]. Across multiple metrics analyzing data fidelity and downstream machine learning task performance, our results show that MPS outperforms classical models, particularly under strict privacy constraints. This work highlights MPS as a promising tool for privacy-aware synthetic data generation. By combining the expressive power of tensor network representations with formal privacy mechanisms, the proposed approach offers an interpretable and scalable alternative for secure data sharing. Its structured design facilitates integration into sensitive domains where both data quality and confidentiality are critical.
@@ I Introduction
Despite its benefits-scalability, accessibility, and flexibility-synthetic data also poses risks. Since it is generated by models trained on real datasets, it may inadvertently leak sensitive information, raising concerns about privacy and confidentiality [5]. Differential Privacy (DP) has emerged as a principled solution to these challenges, offering formal guarantees by injecting calibrated noise into the data generation process. When properly implemented, DP ensures that the inclusion or exclusion of any individual in the dataset does not significantly affect the output.
Concurrently, Tensor Networks (TNs)-a class of models from quantum physics-offer promising tools for high-dimensional data representation. Among them, Matrix Product States (MPS) provide a structured and interpretable parameterization of complex joint distributions, with linear scaling in the number of features. While MPS models have shown success in representing binary data [1], their application to mixed-type synthetic tabular data remains underexplored.
@@ II.1 Deep Generative Models for Tabular Data
The generation of synthetic tabular data has garnered increasing attention with the rise of data-driven applications constrained by privacy, regulatory, or availability concerns. Among the most prominent techniques are Generative Adversarial Networks (GANs) [6] and Variational Autoencoders (VAEs) [7], which have shown strong empirical performance in generating realistic, high-dimensional data. However, their application to tabular domains introduces unique challenges, such as modeling both discrete and continuous variables, managing imbalanced categorical features, and preserving high-fidelity dependencies across features.
CTGAN [2], a GAN-based framework tailored for tabular data, addresses these issues by employing conditional generation and specialized training objectives to avoid mode collapse. It has become a benchmark model due to its ability to capture complex joint distributions. However, CTGAN requires extensive architecture tuning and lacks native support for formal privacy mechanisms.
@@ II.2 Privacy Mechanisms in Generative Modeling
Differential Privacy (DP) [8] provides a principled framework for protecting individual data entries in statistical computations. A randomized mechanism $\mathcal{M}$ satisfies $(\epsilon,\delta)$-DP if for all measurable sets $S$ and for any pair of neighboring datasets $D$ and $D^{\prime}$ differing in one element:
@@
In deep learning, DP is typically achieved by injecting noise into gradients during training, combined with gradient clipping to bound sensitivity [5]. Privacy loss over time is tracked using frameworks such as Rényi Differential Privacy (RDP) [9] and the moment accountant technique.
While DP-GANs and DP-VAEs have been proposed, the inherent instability of training under noise often leads to utility degradation. PrivBayes [10] adopts an alternative approach by training Bayesian networks with explicit DP constraints. It achieves strong privacy guarantees but suffers from scalability limitations in high-dimensional settings.
@@ II.3 Tensor Networks and Born Machines
Tensor Networks (TNs) [11, 12] originated in quantum physics as a way to efficiently represent high-dimensional entangled states. Among these, the Matrix Product State (MPS)-also known as the tensor train-factorizes a high-order tensor into a sequential chain of low-rank tensors, enabling efficient modeling of global correlations with parameters that scale linearly in the number of variables.
@@
In the context of generative modeling, the Born Machine [13] defines the probability of a binary configuration $x$ as the squared amplitude of a quantum wavefunction:
@@
Tensor networks have demonstrated promise in machine learning applications, including supervised classification [14], clustering [15], anomaly detection [16, 17], and solving PDEs via neural networks [18]. However, their use in synthetic data generation for mixed-type tabular datasets remains limited. Moreover, privacy-preserving extensions to TN-based generative models are scarcely explored. A notable exception is Pozas-Kerstjens et al. [19], which investigates gauge-induced privacy within MPS and neural networks. Unlike our work, which integrates privacy directly into the training algorithm using noise injection and formal DP accounting, that approach focuses on structural transformations post-training.
@@ III Problem Formulation
Integrating DP into synthetic data generation introduces a fundamental challenge: maintaining the utility of the synthetic data while safeguarding privacy. To address this, during the MPS model training we integrate stochastic gradient descent (SGD) with gradient clipping and noise injection. These mechanisms are compatible with DP accounting methods, particularly moment accounting based on Rényi Differential Privacy (RDP) [5, 20], allowing us to track cumulative privacy loss and set strict guarantees on the $\epsilon$-differential privacy budget. These techniques ensure that sensitive information is obscured while minimizing performance degradation in downstream tasks.
@@ III.1 Dataset description
To evaluate the performance of our MPS-based synthetic data generator, we use the well-established Adult Income dataset from the 1996 U.S. Census Bureau [21]. This dataset contains 48,842 records with a mix of 6 integer/continuous and 8 categorical features, offering a representative benchmark for mixed-type tabular data.
@@ IV.3 Noise Injection
We adopt the moment accounting method from Abadi et al. [5] to track cumulative privacy loss during SGD. This approach rigorously quantifies the privacy-utility trade-off but requires careful calibration of the noise distribution and its interaction with model parameters. Noise is added at each gradient step, with magnitude calibrated based on dataset sampling ratio and privacy parameters according to:
@@ IV.5.1 Fidelity Metrics
Fidelity scores are computed using the SDMetrics library [22], which provides standardized evaluation tools for synthetic tabular data. Each metric is evaluated over 10 synthetic samples per configuration. These metrics assess the statistical similarity and distance between synthetic and real data columns:
@@ VII Acknowledgments
This work is a collaboration between Bundesdrekerei Group and Multiverse Computing. We would like to acknowledge and thank the technical teams and field engineers both at Bundesdrekerei Group and Multiverse Computing for extremely helpful discussions.

### v2

@@ Abstract
Synthetic data generation is a key technique in modern artificial intelligence, addressing data scarcity, privacy constraints, and the need for diverse datasets in training robust models. In this work, we propose a method for generating privacy preserving high-quality synthetic tabular data using Tensor Networks, specifically Matrix Product States (MPS) [Wang2023].We benchmark the MPS-based generative model against state-of-the-art models such as CTGAN [xu2018synthesizing], VAE, and PrivBayes [golob2024high], focusing on both fidelity and privacy-preserving capabilities. To ensure differential privacy (DP), we integrate noise injection and gradient clipping during training, enabling privacy guarantees via Rényi Differential Privacy accounting [giomi2022unified]. Across multiple metrics analyzing data fidelity and downstream machine learning task performance, our results show that MPS outperforms classical models, particularly under strict privacy constraints. This work highlights MPS as a promising tool for privacy-aware synthetic data generation. By combining the expressive power of tensor network representations with formal privacy mechanisms, the proposed approach offers an interpretable and scalable alternative for secure data sharing. Its structured design facilitates integration into sensitive domains where both data quality and confidentiality are critical.
@@ I Introduction
Despite its benefits-scalability, accessibility, and flexibility-synthetic data also poses risks. Since it is generated by models trained on real datasets, it may inadvertently leak sensitive information, raising concerns about privacy and confidentiality [Abadi_2016]. Differential Privacy (DP) has emerged as a principled solution to these challenges, offering formal guarantees by injecting calibrated noise into the data generation process. When properly implemented, DP ensures that the inclusion or exclusion of any individual in the dataset does not significantly affect the output.
Concurrently, Tensor Networks (TNs)-a class of models from quantum physics-offer promising tools for high-dimensional data representation. Among them, Matrix Product States (MPS) provide a structured and interpretable parameterization of complex joint distributions, with linear scaling in the number of features. While MPS models have shown success in representing binary data [Wang2023], their application to mixed-type synthetic tabular data remains underexplored.
@@ II.1 Deep Generative Models for Tabular Data
The generation of synthetic tabular data has garnered increasing attention with the rise of data-driven applications constrained by privacy, regulatory, or availability concerns. Among the most prominent techniques are Generative Adversarial Networks (GANs) [goodfellow2014generative] and Variational Autoencoders (VAEs) [kingma2013auto], which have shown strong empirical performance in generating realistic, high-dimensional data. However, their application to tabular domains introduces unique challenges, such as modeling both discrete and continuous variables, managing imbalanced categorical features, and preserving high-fidelity dependencies across features.
CTGAN [xu2018synthesizing], a GAN-based framework tailored for tabular data, addresses these issues by employing conditional generation and specialized training objectives to avoid mode collapse. It has become a benchmark model due to its ability to capture complex joint distributions. However, CTGAN requires extensive architecture tuning and lacks native support for formal privacy mechanisms.
@@ II.2 Privacy Mechanisms in Generative Modeling
Differential Privacy (DP) [Dwork2006Calibrating] provides a principled framework for protecting individual data entries in statistical computations. A randomized mechanism $\mathcal{M}$ satisfies $(\epsilon,\delta)$-DP if for all measurable sets $S$ and for any pair of neighboring datasets $D$ and $D^{\prime}$ differing in one element:
@@
In deep learning, DP is typically achieved by injecting noise into gradients during training, combined with gradient clipping to bound sensitivity [Abadi_2016]. Privacy loss over time is tracked using frameworks such as Rényi Differential Privacy (RDP) [Mironov2017RDP] and the moment accountant technique.
While DP-GANs and DP-VAEs have been proposed, the inherent instability of training under noise often leads to utility degradation. PrivBayes [zhang2017privbayes] adopts an alternative approach by training Bayesian networks with explicit DP constraints. It achieves strong privacy guarantees but suffers from scalability limitations in high-dimensional settings.
@@ II.3 Tensor Networks and Born Machines
Tensor Networks (TNs) [orus2014practical, Verstraete_2008] originated in quantum physics as a way to efficiently represent high-dimensional entangled states. Among these, the Matrix Product State (MPS)-also known as the tensor train-factorizes a high-order tensor into a sequential chain of low-rank tensors, enabling efficient modeling of global correlations with parameters that scale linearly in the number of variables.
@@
In the context of generative modeling, the Born Machine [Han2018MPSgen] defines the probability of a binary configuration $x$ as the squared amplitude of a quantum wavefunction:
@@
Tensor networks have demonstrated promise in machine learning applications, including supervised classification [Stoudenmire2016SupervisedLW], clustering [Stoudenmire_2018], anomaly detection [Wang2020, aizpurua2023], and solving PDEs via neural networks [Patel2022]. However, their use in synthetic data generation for mixed-type tabular datasets remains limited. Moreover, privacy-preserving extensions to TN-based generative models are scarcely explored. A notable exception is Pozas-Kerstjens et al. [pozas2024privacy], which investigates gauge-induced privacy within MPS and neural networks. Unlike our work, which integrates privacy directly into the training algorithm using noise injection and formal DP accounting, that approach focuses on structural transformations post-training.
@@ III Problem Formulation
Integrating DP into synthetic data generation introduces a fundamental challenge: maintaining the utility of the synthetic data while safeguarding privacy. To address this, during the MPS model training we integrate stochastic gradient descent (SGD) with gradient clipping and noise injection. These mechanisms are compatible with DP accounting methods, particularly moment accounting based on Rényi Differential Privacy (RDP) [Abadi_2016, Mironov_2017], allowing us to track cumulative privacy loss and set strict guarantees on the $\epsilon$-differential privacy budget. These techniques ensure that sensitive information is obscured while minimizing performance degradation in downstream tasks.
@@ III.1 Dataset description
To evaluate the performance of our MPS-based synthetic data generator, we use the well-established Adult Income dataset from the 1996 U.S. Census Bureau [Dataset]. This dataset contains 48,842 records with a mix of 6 integer/continuous and 8 categorical features, offering a representative benchmark for mixed-type tabular data.
@@ IV.3 Noise Injection
We adopt the moment accounting method from Abadi et al. [Abadi_2016] to track cumulative privacy loss during SGD. This approach rigorously quantifies the privacy-utility trade-off but requires careful calibration of the noise distribution and its interaction with model parameters. Noise is added at each gradient step, with magnitude calibrated based on dataset sampling ratio and privacy parameters according to:
@@ IV.5.1 Fidelity Metrics
Fidelity scores are computed using the SDMetrics library [sdmetrics], which provides standardized evaluation tools for synthetic tabular data. Each metric is evaluated over 10 synthetic samples per configuration. These metrics assess the statistical similarity and distance between synthetic and real data columns:
@@ VII Acknowledgments
This work is a collaboration between Bundesdruckerei Group and Multiverse Computing. We would like to acknowledge and thank the technical teams and field engineers both at Bundesdruckerei Group and Multiverse Computing for extremely helpful discussions.